# Does Positive Reinforcement Work?: A Quasi-Experimental Study of the Effects of Positive Feedback on Reddit

### v1

@@ Abstract.
Social media platform design often incorporates explicit signals of positive feedback. Some moderators provide positive feedback with the goal of positive reinforcement, but are often unsure of their ability to actually influence user behavior. Despite its widespread use and theory touting positive feedback as crucial for user motivation, its effect on recipients is relatively unknown. This paper examines how positive feedback impacts Reddit users and evaluates its differential effects to understand who benefits most from receiving positive feedback. Through a causal inference study of 11M posts across 4 months, we find that users who received positive feedback made more frequent (2% per day) and higher quality (57% higher score; 2% fewer removals per day) posts compared to a set of matched control users. Our findings highlight the need for platforms and communities to expand their perspective on moderation and complement punitive approaches with positive reinforcement strategies.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
“We heard you… awards are back!”– Reddit Product Team [^1] https://www.reddit.com/r/reddit/comments/1css0ws/we_heard_you_awards_are_back/
In September 2023, Reddit removed several positive feedback mechanisms, including awards and gilds, from its interface [^2] https://www.reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to_awards_coins_and/ in an effort to create better ways to empower communities to reward content. Although gilding—i.e., the ability to “gild” posts by donating a Reddit Premium subscription—is no longer a feature on Reddit, the platform brought back other forms of awards in May 20241 based on feedback from its users. These recent changes reflect both the platform’s interest in supporting positive feedback and the users’ desire to continue providing such feedback.
Positive feedback is a key feature of positive reinforcement, a long-studied behavior modification principle in the field of behavioral psychology shown to be effective in offline settings to encourage desired behavior through positive stimuli  (Ferster and Skinner, 1957). In the online context, recent HCI research has revealed that some moderators of online communities consider positive reinforcement a part of their role (Lambert et al., 2024). In particular, the system of rewards utilized in positive reinforcement involves the use of the available signals on the platform. Importantly, most social media platforms include some form of positive feedback (e.g., Reddit upvotes, Facebook likes, etc.). In fact, the presence of positive feedback as an affordance is associated with a community consisting of higher-quality content (Papakyriakopoulos et al., 2023). Furthermore, HCI researchers have established that positive feedback increases a user’s motivation to contribute to a community (Kraut et al., 2011). As a result, positive feedback may be vital both for users to want to participate and for platforms that rely on contributions to be successful. Despite the ubiquity of positive feedback across platforms and the established theories motivating its potential to have meaningful impacts offline and online, there is a lack of empirical evidence supporting the effectiveness of receiving positive feedback on improving user-level outcomes. Learning whether positive feedback can measurably change important outcomes can alter our perspective on moderation and motivate the need for platforms to support proactive approaches to moderation centered around reinforcement in conjunction with preexisting punitive approaches.
@@
To explore distributed moderation through positive feedback, we focus on Reddit gilds and upvotes, two forms of positive feedback available to all community members. Figure 1 provides an example post that received both forms of feedback. Gilds were a beloved (Weatherbed, 2024) but sparsely-used, paid feature while upvotes are free to give out and used in all sub-communities on Reddit. Prior research confirms that moderators utilize both gilds and upvotes as mechanisms for positive reinforcement (Lambert et al., 2024), however both forms of feedback are given anonymously, so we cannot evaluate their efficacy as moderator-specific feedback and instead explore whether they are capable of accomplishing the goals of moderation when given by the community more generally (i.e., distributed).
Figure 1. This is an example post from r/aww that received a gild. The gild is represented as a small gold icon above the post title and score is reported under the post title. This post received a score in the top 25th percentile for its subreddit in the month it was posted.
@@
Reddit post from r/aww that contains a picture of a small cat in a sink. The post received one gild and a score of 10K.
@@ 1.1. Research Questions
The posting quality, posting frequency, and norm-adherence of users are important to communities (Weld et al., 2024), platforms (Kraut et al., 2011), and moderators (Lambert et al., 2024) alike. Thus, we ask whether positive reinforcement through distributed moderation is capable of influencing these important user-level outcomes. More specifically, we pose the following research questions:
@@ 1. RQ1:
      How does receiving a gild or being highly upvoted impact users’ future behavior in a community?
@@ 2. RQ2:
    How do the effects of positive feedback vary between recipients?
@@ 1.2. Summary of Contributions
With these research questions in mind, we employ causal inference methods to estimate the causal effects of receiving positive feedback (i.e., our treatments). We first form a group of treatment posts which received a large quantity of upvotes and/or received a gild. We then form a control group of similar posts from the same communities that did not receive such forms of positive feedback. With these two groups, we perform a difference-in-differences analysis to understand the effect of receiving positive feedback on a user’s future behavior in their community and compare pre- and post-treatment behavior to understand how long these effects persist.
We then explore the differential effects of our treatments through regressions on the individual treatment effects to identify which types of users and communities see stronger effects of receiving positive feedback.
@@
We find that users who receive gilds or high score on a post receive more positive feedback on future posts compared to if they had never received the positive feedback. Highly-upvoted users also experience an increase in their posting rate and a decrease in their rate of removals. Additionally, we observe that users who received stronger treatment (i.e., more gilds or upvotes) experience larger increases in future positive feedback received. Importantly, Reddit newcomers experience increased effects of being highly-upvoted on their future post and removal rates.
@@ 2.1. Principles in Behavioral Psychology
One specific type of reinforcing feedback that has been studied is praise. Kazdin (1978) find that praising desired behavior increases the likelihood that the behavior will be repeated. On Reddit, praise happens through upvoting and awarding signals. Prior work found that some moderators specifically provide positive feedback like gilds and upvotes to reinforce behavior and others specifically aimed to praise the author (Lambert et al., 2024).
Additional research has identified another type of reinforcement: vicarious. The idea of vicarious reinforcement is that bystanders of positive reinforcement may be vicariously influenced to repeat the behavior being reinforced (Bandura and Barab, 1971; Kazdin, 1973). Jhaver et al. (2024) explore the effect of exposure to post-removal explanations on bystanders, essentially vicarious punishment. Though they found that the exposure did not encourage bystanders to learn the norms of the community, prior work in in social computing shows that vicarious reinforcement through highlighting examples of norm-adhering behavior in a community may help with further norm-adherence (Kiesler et al., 2012). Seering et al. (2017) showed that users on Twitch imitated observed behaviors, especially from high status users. Furthermore, some users believe providing positive feedback can increase a post’s visibility in the feed (Lambert et al., 2024), effectively highlighting the content for the rest of the community. This exposure to posts approved of by the community may enable community norm acquisition by newcomers and other bystanders through vicarious reinforcement.
Miltenberger (2016) further explored the factors that may affect the effectiveness of positive reinforcement strategies in practice and presents three notable factors: contingency, satiation, and immediacy. Contingency states that the effect of a positive stimulus will be strongest when it is only presented in response to the behavior being reinforced and at no other time. Satiation is the idea that a positive stimulus will have a diminished effect on the recipient if they have been exposed to the stimulus too much. Wang and Diakopoulos (2021) demonstrate the idea of satiation in their work by finding that receiving a New York Times Pick badge, a form of positive reinforcement, has the largest effect on first-time recipients’ future comment quality and has diminishing effects for more frequent recipients. Finally, the concept of immediacy indicates that receiving the reinforcing stimulus more quickly will be more effective. Choi et al. (2024) explored the idea of immediacy in the context of YouTube creator hearts and found that videos received more engagement if creator hearts are given sooner after the video was published.
@@ 2.2. Distributed Moderation
Reddit’s moderation system is most often categorized as an example of centralized moderation. While each sub-community is able to create its own set of rules, they are enforced with centralized teams of volunteer moderators. These moderation teams are crucial to user safety, but the users themselves play a similarly important role in moderation. Reddit enables users to give various forms of feedback for contributions, such as upvotes, downvotes, and awards. These signals are then taken into consideration for the sorting of community feeds. In this way, users have the collective power to highlight content they approve of and hide content they do not. In practice, this is an example of distributed moderation, a system through which members of a community are responsible for making moderation decisions.
There is extensive research on Reddit moderation, including identifying abusive behavior (Chandrasekharan et al., 2017b; Habib et al., 2019; Urbaniak et al., 2022; Almerekhi et al., 2020; Bagga et al., 2021), evaluating the efficacy of moderation strategies (Chandrasekharan et al., 2022, 2017a; Srinivasan et al., 2019; Jhaver et al., 2019b; Matias, 2019), and characterizing moderator experiences (Schaffner et al., 2024; Kiene and Hill, 2020). However, prior work does not often explore the power of Reddit’s distributed moderation through user feedback, though some Reddit moderators themselves have stated that users should utilize upvotes and downvotes on the platform to self-regulate, demonstrating support for distributed moderation from active participants in Reddit’s centralized moderation (Lambert et al., 2024).
Prior work into distributed moderation often focuses on Slashdot. Lampe and Resnick (2004) and Lampe et al. (2014) show the power distributed moderation on Slashdot has to identify high and low quality contributions, while also reporting some of the challenges associated with the approach. Jiang et al. (2023) and Grimmelmann (2015) similarly identify trade-offs between various moderation practices, including comparisons of centralized and distributed moderation. Jiang et al. (2023) deem distributed moderation a more democratic process as it better reflects community-wide opinions.
@@ 2.3. Guiding User Behavior Through Interface Signals
Platform designers have the power to determine what users can do in their communities by choosing to include or exclude various interface elements. For example, some platforms utilize badges (Anderson et al., 2013), gamification mechanisms (Papoutsoglou et al., 2020; Hamari et al., 2014; Sailer et al., 2017), or reputation systems (Resnick et al., 2000, 2006; Adler and De Alfaro, 2007), which have all been shown to encourage participation. Prior work has also shown that users can be encouraged to make higher-quality contributions through the use of specific highlighting mechanisms (Wang and Diakopoulos, 2021). Similarly, posts on Reddit receive one upvote by default, an instance of example-setting that may encourage other users to follow suit. Finally, other research has shown that prominently displaying community rules and guidelines encourages adherence, especially from newcomers (Matias, 2019). This practice is used across Reddit and Twitch. [^3] https://safety.twitch.tv/s/article/Chat-Tools
Social computing researchers have developed new tools and signals on platforms to encourage certain behaviors. Im et al. (2020), for example, introduced a signal to the Twitter interface to inform users about an account’s history with toxicity and misinformation. Chang et al. (2022) introduce an intervention for Reddit users engaging with inflammatory content to encourage more reflection about whether they should contribute to uncivil conversations. Similarly, Yen et al. (2023) build a tool to visualize the amount of negativity in the chatroom of live-streaming platforms to encourage viewers to proactively moderate the chatroom space by contributing prosocially.
@@
Similarly, platforms can consider social translucence (Gilbert, 2012; Erickson and Kellogg, 2000), the concept of emphasizing the visibility of social information in platform design, when making decisions. For example, X (formerly Twitter) has recently changed the visibility of users’ liked posts so others can no longer see who liked another user’s post, nor can users see all the posts other users have liked. [^4] https://x.com/XEng/status/1800634371906380067 Similarly, Instagram now provides the option to make the like count on posts private. [^5] https://about.instagram.com/blog/announcements/giving-people-more-control Ostensibly these changes are motivated by user privacy or maintaining user well-being, but there is disagreement in prior work regarding whether like counts on Instagram are responsible for negative effects on user well-being (Wallace and Buil, 2021). Additionally, there are negative side effects of these adjustments on social translucence because they mask vital network information in a way that may hinder interactions between users.
Certain interface features, however, potentially enable users to influence their own communities. Reddit upvotes, for example, are a widespread mechanism used across the platform and are controlled entirely by Reddit users. There has been a lot of research in understanding the role of upvotes on Reddit (Morrison and Hayes, [n. d.]), including in spaces like advice-giving (Lu, 2024) and political communities (Papakyriakopoulos et al., 2023; Carman et al., 2018). Prior work has shown that interfaces that allow upvotes typically have higher-quality content (Papakyriakopoulos et al., 2023). Additionally, Carman et al. (2018) manipulated the number of upvotes given to posts in political and non-political subreddits and found that posts with artificially-inflated upvote counts experienced increases in their final upvote and reply counts, demonstrating the power of the upvoting mechanism to increase engagement and positive feedback. We extend this research to understand the specific influence positive feedback has on the recipients themselves.

### v2

@@ Abstract.
Social media platform design often incorporates explicit signals of positive feedback. Some moderators provide positive feedback with the goal of positive reinforcement, but are often unsure of their ability to actually influence user behavior. Despite its widespread use and theory touting positive feedback as crucial for user motivation, its effect on recipients is relatively unknown. This paper examines how positive feedback impacts Reddit users and evaluates its differential effects to understand who benefits most from receiving positive feedback. Through a causal inference study of 11M posts across 4 months, we find that users who received positive feedback made more frequent (2% per day) and higher quality (57% higher score; 2% fewer removals per day) posts compared to a set of matched control users. Our findings highlight the need for platforms, communities, and moderators to expand their perspective on moderation and complement punitive approaches with positive reinforcement strategies to foster desirable behavior online.
@@
journalyear: 2025copyright: noneconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713830isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
“We heard you… awards are back!”– Reddit Product Team [^1] https://reddit.com/r/reddit/comments/1css0ws/we_heard_you_awards_are_back/
In September 2023, Reddit removed several positive feedback mechanisms, including awards and Reddit gold, from its interface [^2] https://reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to_awards_coins_and/ in an effort to create better ways to empower communities to reward content. Although gilding—i.e., the ability to “gild” posts by donating a Reddit Premium subscription—is no longer a feature on Reddit, the platform brought back other forms of awards in May 20241 based on feedback from its users. These recent changes reflect both the platform’s interest in supporting positive feedback and the users’ desire to continue providing such feedback.
Positive feedback is a key feature of positive reinforcement, a long-studied behavior modification principle in the field of behavioral psychology shown to be effective in offline settings to encourage desired behavior through positive stimuli  (Ferster and Skinner, 1957). In the online context, recent HCI research has revealed that some moderators of online communities consider positive reinforcement a part of their role (Lambert et al., 2024a). In particular, the system of rewards utilized in positive reinforcement involves the use of the available signals on the platform. Importantly, most social media platforms include some form of positive feedback (e.g., Reddit upvotes, Facebook likes, etc.). In fact, the presence of positive feedback as an affordance is associated with a community consisting of higher-quality content (Papakyriakopoulos et al., 2023). Furthermore, HCI researchers have established that positive feedback increases a user’s motivation to contribute to a community (Kraut et al., 2011). As a result, positive feedback may be vital both for users to want to participate and for platforms that rely on contributions to be successful. Despite the ubiquity of positive feedback across platforms and the established theories motivating its potential to have meaningful impacts offline and online, there is a lack of empirical evidence supporting the effectiveness of receiving positive feedback on improving user-level outcomes. Learning whether positive feedback can measurably change important outcomes can alter our perspective on moderation and motivate the need for platforms to support proactive approaches to moderation centered around reinforcement in conjunction with preexisting punitive approaches.
@@
To explore distributed moderation through positive feedback, we focus on Reddit gold and upvotes, two forms of positive feedback available to all community members. Figure 1 provides an example post that received both forms of feedback. Gold was a beloved (Weatherbed, 2024) but sparsely-used, paid feature while upvotes are free to give out and used in all sub-communities on Reddit. Prior research confirms that moderators utilize both gold and upvotes as mechanisms for positive reinforcement (Lambert et al., 2024a), however both forms of feedback are given anonymously, so we cannot evaluate their efficacy as moderator-specific feedback and instead explore whether they are capable of accomplishing the goals of moderation when given by the community more generally (i.e., distributed).
Figure 1. This is an example post from r/aww that received gold, represented by a small gold icon above the title. The post’s score is reported under its title. This post’s score was in the top 25th percentile for r/aww in the month it was posted.
@@
Reddit post from r/aww that contains a picture of a small cat in a sink. The post received one gold and a score of 10K.
@@ 1.1. Research Questions
The posting quality, posting frequency, and norm-adherence of users are important to communities (Weld et al., 2024), platforms (Kraut et al., 2011), and moderators (Lambert et al., 2024a) alike. Thus, we ask whether positive reinforcement through distributed moderation is capable of influencing these important user-level outcomes. More specifically, we pose the following research questions:
@@ 1. RQ1:
      How does receiving gold or being highly upvoted impact users’ future behavior in a community?
@@ 2. RQ2:
    How do the effects of positive feedback vary across different types of recipients?
@@ 1.2. Summary of Contributions
With these research questions in mind, we employ causal inference methods to estimate the causal effects of receiving positive feedback (i.e., our treatments). In particular, we draw on the potential outcomes framework (Rubin, 2005; Imbens and Rubin, 2015) to conduct stratified matching followed by difference-in-difference analyses. To do so, we first form a group of treatment posts which received a large quantity of upvotes and/or received gold. We then form a control group of similar posts from the same communities that did not receive such forms of positive feedback. With these two groups, we perform stratified matching to match users with similar propensities of being treated. Then, we conduct a difference-in-differences analysis to understand the effect of receiving positive feedback on a user’s future behavior in their community and compare pre- and post-treatment behavior to understand how long these effects persist.
We then explore the differential effects of our treatments through regressions on the individual treatment effects to identify which types of users and communities experience stronger effects of receiving positive feedback.
@@
We find that users who receive gold or high score on a post receive more positive feedback on future posts compared to if they had never received the positive feedback. Highly-upvoted users also experience an increase in their posting rate and a decrease in their rate of removals. Additionally, we observe that users who received stronger treatment (i.e., more gold or upvotes) experience larger increases in future positive feedback received. Importantly, Reddit newcomers experience larger increases in posting frequency and larger decreases in removal rates after being highly upvoted.
@@ 2.1. Principles in Behavioral Psychology
One specific type of reinforcing feedback that has been studied is praise. Kazdin (1978) find that praising desired behavior increases the likelihood that the behavior will be repeated. On Reddit, praise happens through upvoting and awarding signals. Prior work found that some moderators specifically provide positive feedback like gold and upvotes to reinforce behavior and others specifically aimed to praise the author (Lambert et al., 2024a).
Additional research has identified another type of reinforcement: vicarious. The idea of vicarious reinforcement is that bystanders of positive reinforcement may be vicariously influenced to repeat the behavior being reinforced (Bandura and Barab, 1971; Kazdin, 1973). Jhaver et al. (2024) explore the effect of exposure to post-removal explanations on bystanders, essentially vicarious punishment. Though they found that the exposure did not encourage bystanders to learn the norms of the community, prior work in in social computing shows that vicarious reinforcement through highlighting examples of norm-adhering behavior in a community may help with further norm-adherence (Kiesler et al., 2012). Seering et al. (2017) showed that users on Twitch imitated observed behaviors, especially from high status users. Furthermore, some users believe providing positive feedback can increase a post’s visibility in the feed (Lambert et al., 2024a), effectively highlighting the content for the rest of the community. This exposure to posts approved of by the community may enable community norm acquisition by newcomers and other bystanders through vicarious reinforcement.
Miltenberger (2016) further explored the factors that may impact the effectiveness of positive reinforcement strategies in practice and presents three notable factors: contingency, satiation, and immediacy. Contingency states that a positive stimulus will have the strongest effect when it is only presented in response to the behavior being reinforced and at no other time. Satiation is the idea that a positive stimulus will have a diminished effect on the recipient if they have been too exposed to the stimulus. Wang and Diakopoulos (2021) demonstrate the idea of satiation in their work which reveals that receiving a New York Times Pick badge, a form of positive reinforcement, has the largest effect on first-time recipients’ future comment quality and diminishing effects for more frequent recipients. Finally, the concept of immediacy indicates that the reinforcing stimulus will be more effective when received more quickly. Choi et al. (2025) explored the idea of immediacy in the context of YouTube creator hearts and found that videos received more engagement when creator hearts are given sooner after the video was published.
@@ 2.2. Distributed Moderation
Reddit’s moderation system exhibits qualities of both centralized and distributed moderation systems. There are some platform-level moderation structures (e.g., admins [^3] https://redditinc.com/policies/content-policy), but sub-communities create their own sets of rules (Reddy and Chandrasekharan, 2023) and most moderation falls to volunteer moderation teams within each community. While these volunteers themselves are Reddit users, they still form centralized moderation structures within communities. These moderation teams are crucial to user safety, but the non-moderating users play a similarly important role in moderation. Reddit enables users to give various forms of feedback to contributions, such as upvotes, downvotes, and awards. These signals are then taken into consideration for the sorting of community feeds. In this way, users have the collective power to highlight content they approve of and hide content they do not. In practice, this is an example of distributed moderation, a system through which members of a community are responsible for making moderation decisions.
There is extensive research on Reddit moderation, including identifying abusive behavior (Chandrasekharan et al., 2017b; Habib et al., 2019; Urbaniak et al., 2022; Almerekhi et al., 2020; Bagga et al., 2021), evaluating the efficacy of moderation strategies (Chandrasekharan et al., 2022, 2017a; Srinivasan et al., 2019; Jhaver et al., 2019b; Matias, 2019), and characterizing moderator experiences (Schaffner et al., 2024; Kiene and Hill, 2020). However, prior work does not often explore the power of Reddit’s distributed moderation through user feedback, though some Reddit moderators themselves have stated that users should utilize upvotes and downvotes on the platform to self-regulate, demonstrating support for distributed moderation from active participants in Reddit’s centralized moderation (Lambert et al., 2024a).
Prior work into distributed moderation often focuses on the platform Slashdot. Lampe and Johnston (2005) specifically highlight the effect of positive and negative feedback given to an author’s first contribution on their future participation on Slashdot. Additionally, Lampe and Resnick (2004) and Lampe et al. (2014) show the power distributed moderation on Slashdot has to identify high and low quality contributions, while also reporting some of the challenges associated with the approach. Jiang et al. (2023) and Grimmelmann (2015) similarly identify trade-offs between various moderation practices, including comparisons of centralized and distributed moderation. Jiang et al. (2023) deem distributed moderation a more democratic process as it better reflects community-wide opinions.
@@ 2.3. Guiding User Behavior Through Interface Signals
Platform designers have the power to determine what users can do in their communities by choosing to include or exclude various interface elements. For example, some platforms utilize badges (Anderson et al., 2013), gamification mechanisms (Papoutsoglou et al., 2020; Hamari et al., 2014; Sailer et al., 2017), or reputation systems (Resnick et al., 2000, 2006; Adler and De Alfaro, 2007), which have all been shown to encourage participation. Prior work has also shown that users can be encouraged to make higher-quality contributions through the use of specific highlighting mechanisms (Wang and Diakopoulos, 2021). Similarly, posts on Reddit receive one upvote by default, an instance of example-setting that may encourage other users to follow suit. Upvotes and other explicit signals of positive feedback on Reddit are examples of formal feedback, which may be more effective at encouraging norm awareness and abidance than informal forms of feedback (Kiesler et al., 2012). Finally, other research has shown that prominently displaying community rules and guidelines encourages adherence, especially from newcomers (Matias, 2019). This practice is used across Reddit and Twitch. [^4] https://safety.twitch.tv/s/article/Chat-Tools
Social computing researchers have developed new tools and signals on platforms to encourage certain behaviors. Im et al. (2020), for example, introduced a signal to the Twitter interface to inform users about an account’s history with toxicity and misinformation. Chang et al. (2022) introduce an intervention for Reddit users engaging with inflammatory content to encourage more reflection about whether they should contribute to uncivil conversations. Similarly, Yen et al. (2023) build a tool to visualize the amount of negativity in live-streaming chatrooms to encourage viewers to proactively moderate the space by contributing prosocially.
@@
Similarly, platforms can consider social translucence (Gilbert, 2012; Erickson and Kellogg, 2000), the concept of emphasizing the visibility of social information in platform design, when making decisions. For example, X (formerly Twitter) has recently changed the visibility of users’ liked posts so others can no longer see who liked another user’s post, nor can users see the posts other users have liked. [^5] https://x.com/XEng/status/1800634371906380067 Similarly, Instagram now provides the option to make the like count on posts private. [^6] https://about.instagram.com/blog/announcements/giving-people-more-control Ostensibly these changes are motivated by user privacy or maintaining user well-being, but there is disagreement in prior work regarding whether like counts on Instagram are responsible for negative effects on user well-being (Wallace and Buil, 2021). Additionally, there are negative side effects of these adjustments on social translucence because they mask vital network information in a way that may hinder interactions between users.
However, certain interface features can enable users to influence their communities. Reddit upvotes, for example, are a widespread mechanism used across the platform and are controlled entirely by Reddit users. There has been prior research in understanding the role of upvotes on Reddit (Morrison and Hayes, 2013; Lambert et al., 2024b), including in spaces like advice-giving (Lu, 2024) and political communities (Papakyriakopoulos et al., 2023; Carman et al., 2018). Prior work has shown that interfaces that allow upvotes typically have higher-quality content (Papakyriakopoulos et al., 2023). Additionally, Carman et al. (2018) manipulated the number of upvotes given to posts in political and non-political subreddits and found that posts with artificially-inflated upvote counts experienced increases in their final upvote and reply counts, demonstrating the power of upvoting to increase engagement and positive feedback. We extend this research to understand the specific influence positive feedback has on the recipients themselves.